実用英語演習 III (その 1) 2 単位 3 年 (前期)

実用英語演習 III (その 1)  座喜純 · 准教授/ 大学院ソシオ・アーツ・アンド・サイエンス研究部

【授業目的】This class formed to give students exposure to and practice with the English writing or speech-making process from collecting ideas to the final version. Students will learn how to build paragraphs and outlines, and proof-reading for English essays or speech presentations.

【授業概要】 English composition and speech presentation 【キーワード】 essay writing， speech presentation， writing in English， speaking

in English 【先行科目】『実用英語演習 I(その 1)』 (1.0)，『実用英語演習 I その 1)』 (1.0) 【関連科目】『実用英語演習 II(その 1)』 (1.0)，『実用英語演習 II(その 2)』 (1.0)，

『実用英語演習 · 総論』 (1.0) 【履修上の注意】 4 つのクラス (水曜 7-8/スティーブンス担当，水曜 3-4 フェネリー

担当，金曜 5-6/早内ジュディス担当，金曜 7-8/マーシェソ担当) のうち， 1 クラスを履修します．履修登録前に希望を聞いた上でクラス指定をします．掲示に注意しておいてください．なお，この授業は学生個人では履修登録できません．

【到達目標】 By the end of this course, students should be able to write and present

competently on almost any topic. 【授業計画】

1. General Class Guidance

2. Guidance for Each Class

3. Introduction: Process Writing

4. Unit 1: Pre-Writing: Getting Ready to Write (1)

5. Unit 1: Pre-Writing: Getting Ready to Write (2)

6. Unit 2: The Structure of a Paragraph (1)

7. Unit 2: The Structure of a Paragraph (2)

8. Unit 3: The Development of a Paragraph (1)

9. Unit 3: The Development of a Paragraph (2)

10. Unit 4: Descriptive and Process Paragraphs (1)

11. Unit 4: Descriptive and Process Paragraphs (2)

12. Unit 5: Opinion Paragraphs (1)

13. Unit 5: Opinion Paragraphs (2)

14. Unit 6: Comparison/ Contrast Paragraph (1)

15. Unit 6: Comparison/ Contrast Paragraph (2)

16. Final Writing or Short Presentation 【成績評価】 Final essay or speech presentation is an important component of this course, but weekly attendance, participation to class work, and homework are even more important. Tentative evaluation scheme: Weekly class participation, assignment, and homework (70%), Final essay or presentation (30%).

【再試験】 Yes, if you have attended more than 2/3 of the class weeks and if your

score is 50% and over. 【教科書】 Dorothy E. Zemach & Lisa A. Rumisek. Success with College Writing:

From Paragraph to Essay. Macmillan Languagehouse, 2003. 【参考書】 [参考資料] 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218671 【連絡先】

⇒ 座喜 . (オフィスアワー: jzaki@ias.tokushima-u.ac.jp)

【備考】 All the four classes are taught by native speakers of English.

実用英語演習 III (その 1) 2 units 3rd-year(1st semester)

Jun Zaki · Associate Professor / Institute of Socio-Arts and Sciences

Target⟩ This class formed to give students exposure to and practice with the English writing or speech-making process from collecting ideas to the final version. Students will learn how to build paragraphs and outlines, and proof-reading for English essays or speech presentations.

Outline⟩ English composition and speech presentation �������⟩ essay writing, speech presentation, writing in English, speaking in

English Fundamental Lecture⟩ “実用英語演習 I(その 1)”(1.0), “実用英語演習 I その 1)”

(1.0)

Relational Lecture⟩ “実用英語演習 II(その 1)”(1.0), “実用英語演習 II(その 2)”

(1.0), “実用英語演習 · 総論”(1.0)

Notice⟩ 1. Students must choose one of the four class periods of this course for each semester. 2. Students will be expected to speak and write only in English at class. They will be expected to work not only in class but also out of class; there will be homework and they are expected to do it even if they are absent from class. 3. Following class schedule is tentative and only shows general contents of the textbook used in this course. It may be modified due to the size of each class or other factors.

Goal⟩ By the end of this course, students should be able to write and present

competently on almost any topic.

Schedule⟩

1. General Class Guidance

2. Guidance for Each Class

3. Introduction: Process Writing

4. Unit 1: Pre-Writing: Getting Ready to Write (1)

5. Unit 1: Pre-Writing: Getting Ready to Write (2)

6. Unit 2: The Structure of a Paragraph (1)

7. Unit 2: The Structure of a Paragraph (2)

8. Unit 3: The Development of a Paragraph (1)

9. Unit 3: The Development of a Paragraph (2)

10. Unit 4: Descriptive and Process Paragraphs (1)

11. Unit 4: Descriptive and Process Paragraphs (2)

12. Unit 5: Opinion Paragraphs (1)

13. Unit 5: Opinion Paragraphs (2)

14. Unit 6: Comparison/ Contrast Paragraph (1)

15. Unit 6: Comparison/ Contrast Paragraph (2)

16. Final Writing or Short Presentation

Evaluation Criteria⟩ Final essay or speech presentation is an important component of this course, but weekly attendance, participation to class work, and homework are even more important. Tentative evaluation scheme: Weekly class participation, assignment, and homework (70%), Final essay or presentation (30%).

Re-evaluation⟩ Yes, if you have attended more than 2/3 of the class weeks and if

your score is 50% and over. Textbook⟩ Dorothy E. Zemach & Lisa A. Rumisek. Success with College Writing:

From Paragraph to Essay. Macmillan Languagehouse, 2003. Reference⟩ [参考資料] Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=218671 Contact⟩

⇒ Zaki . (Office Hour: jzaki@ias.tokushima-u.ac.jp)

Note⟩ All the four classes are taught by native speakers of English.